# Iterative Corpus Refinement for Materials Property Prediction Based on Scientific Texts

### v1

@@ Abstract
Keywords:
text mining materials informatics corpus refinement word embeddings word2vec electrocatalysis
1 Introduction
@@
Despite its potential, using the scientific literature effectively presents challenges. Not all text sources are equally relevant, and including too many irrelevant or redundant sources can reduce the predictive quality of the models. The increasing volume of scientific literature, including AI-generated content, makes this issue even more pressing. Many AI-generated texts are repetitive or inaccurate, introducing noise into the dataset, and if repetetively used to retrain the models eventually leading to their collapse [20, 21]. Effective methods are therefore needed to filter out ‘low-quality content’ to ensure that models are trained on the most meaningful information.
@@ 2 Experiments
2.1 Corpus Collection and Preprocessing
Fig. 1 shows a schematic overview of our methodology. First, we collect a relevant corpus of documents using the PaperCollector module in MatNexus [28]. As source for the documents we use the APIs of Scopus and and limit ourselves to use open access-only publications up to including the year 2023 which results in 6506 papers. The retrieved abstracts and metadata are stored in a structured CSV file. Text preprocessing is performed using the TextProcessor module, which removes licensing statements, filters common English stopwords, and retains domain-specific terms, such as chemical element symbols. Tokenization is applied to prepare the text for embedding generation.
2.2 Embedding Generation and Greedy Selection
The initial document embeddings are generated using a Doc2Vec model trained on the full corpus containing 6506 documents. The embeddings capture the semantic content of each document, mapping the tokens into a 200-dimensional vector space. A greedy selection algorithm is then applied to iteratively refine the subset of documents for analysis.
@@ 2.2 Embedding Generation and Greedy Selection
Iterative Expansion: Subsequent documents are selected by identifying the farthest (in cosine distance) from all previously selected documents. It is a classic greedy farthest point sampling except the initial point is selected always the central document in embedding space instead of a random starting point. To improve computational efficiency and focus on principal directions of variance, the 200-dimensional Doc2Vec embeddings are reduced to two dimensions using principal component analysis (PCA) before distance calculations. The selection process continues until a predetermined batch size of 50 additional documents is reached. This process results in an ordered list of the most different documents in batches of 50 and constitutes the basis for subsequent analysis and optimization.
2.3 Iterative Word2Vec Training and Corpus Refinement
A selected subset is then used to train a Word2Vec [14, 8] model via the Vec-Generator module of MatNexus [28]. The skip-gram architecture is used with a vector size of 200, a window size of 5, and hierarchical softmax.
@@ 2.3 Iterative Word2Vec Training and Corpus Refinement
2.4 Pareto Optimization for Candidate Selection
@@
  • For hydrogen evolution (HER) and oxygen reduction (ORR), the objective was to maximize similarity to conductivity while minimizing similarity to dielectric.
  • For oxygen evolution (OER), the objective was to maximize similarity to dielectric while minimizing similarity to conductivity.
@@
3 Results
3.1 Centroid Convergence Over Iterations
@@
3.2 Pareto Analysis with Full Corpus vs. Selected Subset
@@
3.3 Current Density Measurements and Statistical Overview
@@
4 Discussion
Our iterative selection approach consistently results in models that match or exceed the performance of the model built using the full corpus. For ORR and HER systems, the method identifies more negative (better) current densities, indicating higher catalytic performance. For OER, the selected-corpus model suggests the same high-performing compositions as the full corpus model. These results confirm that reducing the training set -— when done strategically —- does not compromise, and even improves the capability to predict high-performance compositions.
4.1 Advantages of Iterative Corpus Selection
@@
4.2 Practical Impact on Electrocatalyst Screening
@@
4.3 Mitigating the Growing Volume of AI-Generated Text
@@
4.4 Limitations and Potential Extensions
@@
  • Hybrid Models: Integrate advanced NLP methods, such as transformer-based encoders (e.g., BERT or a materials-specific derivative MatBERT [24]), to complement the iterative selection process. These models may provide additional layers of semantic evaluation or help identify nuanced relationships that Word2Vec with fixed representations might miss.

### v2

@@ Abstract
preprint: APS/123-QED
I Introduction
@@
Despite its potential, using the scientific literature effectively presents challenges. Not all text sources are equally relevant, and including too many irrelevant or redundant sources can reduce the predictive quality of the models. The increasing volume of scientific literature, including artificial intelligence (AI)-generated content, makes this issue even more pressing. Many AI-generated texts are repetitive or inaccurate, introducing noise into the dataset, and if repetetively used to retrain the models eventually leading to their collapse [20, 21]. Effective methods are therefore needed to filter out ‘low-quality content’ to ensure that models are trained on the most meaningful information.
@@ II Experiments
II.1 Corpus Collection and Preprocessing
Fig. 1 shows a schematic overview of our methodology. First, we collect a relevant corpus of documents using the PaperCollector module in MatNexus [25]. As source for the documents we use the application programming interfaces (APIs) of Scopus and and limit ourselves to use open access-only publications up to including the year 2023 which results in 6506 papers. The retrieved abstracts and metadata are stored in a structured comma-separated values (CSV) file. Text preprocessing is performed using the TextProcessor module, which removes licensing statements, filters common English stopwords, and retains domain-specific terms, such as chemical element symbols. Tokenization is applied to prepare the text for embedding generation.
II.2 Embedding Generation and Greedy Selection
The initial document embeddings are generated using a Doc2Vec model trained on the full corpus containing 6506 documents. The embeddings capture the semantic content of each document, mapping the tokens into a 200-dimensional vector space, which offers a balanced trade-off between computational efficiency and representational capacity for capturing domain-specific linguistic nuances. A greedy selection algorithm is then applied to iteratively refine the subset of documents for analysis.
@@ II.2 Embedding Generation and Greedy Selection
Iterative Expansion: Subsequent documents are selected by identifying the farthest (in cosine distance) from all previously selected documents. It is a classic greedy farthest point sampling except the initial point is selected always the central document in embedding space instead of a random starting point. To improve computational efficiency and focus on principal directions of variance, the 200-dimensional Doc2Vec embeddings are reduced to two dimensions using principal component analysis (PCA) before distance calculations. The selection process continues until a predetermined batch size of 50 additional documents is reached, which serves as a heuristic balance between introducing sufficient diversity in each iteration and maintaining computational manageability. This process results in an ordered list of the most different documents in batches of 50 and constitutes the basis for subsequent analysis and optimization.
A selected subset is then used to train a Word2Vec [18, 23] model via the Vec-Generator module of MatNexus [25]. The skip-gram architecture is used with a vector size of 200, a window size of 5, and hierarchical softmax. The vector size of 200 provides sufficient capacity to capture nuanced relationships between terms while maintaining training efficiency. A window size of 5 is used to reflect a moderate contextual range, allowing the model to learn meaningful co-occurrence patterns without overextending semantic connections. Hierarchical softmax is chosen for its effectiveness in handling large vocabularies with improved training speed over full softmax.
@@
II.3 Pareto Optimization for Candidate Selection
@@
  • For HER and ORR, the objective was to maximize similarity to conductivity while minimizing similarity to dielectric.
  • For oxygen evolution OER, the objective was to maximize similarity to dielectric while minimizing similarity to conductivity.
@@
III Results
III.1 Centroid Convergence Over Iterations
@@
III.2 Pareto Analysis with Full Corpus vs. Selected Subset
@@
III.3 Current Density Measurements and Statistical Overview
@@
IV Discussion
Our iterative selection approach consistently results in models that match or exceed the performance of the model built using the full corpus. For ORR and HER systems, the method identifies more negative (better) current densities, indicating higher catalytic performance. For OER, the selected-corpus model suggests the same high-performing compositions as the full corpus model. These results confirm that reducing the training set—when done strategically—does not compromise, and even improves the capability to predict high-performance compositions.
IV.1 Advantages of Iterative Corpus Selection
@@
IV.2 Practical Impact on Electrocatalyst Screening
@@
IV.3 Mitigating the Growing Volume of AI-Generated Text
@@
IV.4 Limitations and Potential Extensions
@@
  • Hybrid Models: Integrate advanced NLP methods, such as transformer-based encoders (e.g., MatBERT [31]), to complement the iterative selection process. These models may provide additional layers of semantic evaluation or help identify nuanced relationships that Word2Vec with fixed representations might miss.